• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Cole

    Log entry time 16:04:34 on April16,2001

    Entry number 60654

    keyword=Day Shift Summary

    8:00 - Day shift begins in controlled access.

    Cosmic tests were done to ensure the proper function of the spectrometers. There was a problem with the left -HRS Scintillator S1 (SPEC_S1.ADC_L_C4). Things seen to be working now.

    There was no beam at all during the day shift; therefore; no data was taken. Hall A remained in controlled access the entire shift. The survey for the spectrometers was done and the magnet polarities were set. Also, the Quads were cycled.